OpenStudy (anonymous):

I have to find the inverse function on these and Im so confused - how do I do them? ease explain: f^-1(6)=-8 find f(-8)

OpenStudy (anonymous):

I have steps 1 - write in y=f(x) 2- solve the equation in step 1 for x in terms of y 3 - interchange x and y in the equation in step 2

OpenStudy (jdoe0001):

\(\bf f(\textit{some domain}) = \textit{some range}\\ \quad \\ f^{-1}(\textit{some range}) = \textit{some domain}\\ \quad \\ f^{-1}(6) = -8\\ \quad \\ f(8) = \square ?\)

OpenStudy (anonymous):

-8?

OpenStudy (jdoe0001):

:)

OpenStudy (jdoe0001):

the range of the function, is the domain of its inverse the domain of its inverse, is the range of the functoin

OpenStudy (anonymous):

what happens to the 6?

OpenStudy (jdoe0001):

ohh smokes.. I ... right you said .hmm heheh -8 well no :| anyhow.. the idea is that

OpenStudy (anonymous):

its not -8?

OpenStudy (jdoe0001):

say... lemme give you an example \(\bf f(x)= 2x+3\qquad \qquad f^{-1}(x) = \cfrac{x-3}{2}\\ \quad \\ f(\color{red}{5}) = 2(5)+3\implies \color{blue}{13}\qquad \qquad f^{-1}(\color{blue}{13})=\cfrac{(13)-3}{2}\implies \cfrac{10}{2}\implies \color{red}{5}\)

OpenStudy (jdoe0001):

so \(\bf f(\textit{some domain}) = \textit{some range}\\ \quad \\ f^{-1}(\textit{some range}) = \textit{some domain}\\ \quad \\ f^{-1}(6) = -8\\ \quad \\ f(8) = \square ?\)

OpenStudy (anonymous):

6?

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

really?! this makes my head hurt. ... so that doesn't change .... I don't get it =(

OpenStudy (jdoe0001):

if you give a value to "x" in the function, any value then "y" will become, whatever the function spits out if you give whatever the function spit out, as a value to "x" IN THE INVERSE the inverse function will spit out, the original value given to the ORIGINAL function

OpenStudy (anonymous):

ohhhh! f^-1 is 6 and when y=f(x) the 6 would go where the (x) is?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

yay got it! . I hope! thank you

OpenStudy (jdoe0001):

yw